1) Consider the following reaction and associated equilibrium constant:
aA(g)⇌bB(g), Kc = 2.7

Part A
Find the equilibrium concentrations of A and B for a=1 and b=1. Assume that the initial concentration of A is 1.0 M and that no B is present at the beginning of the reaction.

Part B
Find the equilibrium concentrations of A and B for a=2 and b=2. Assume that the initial concentration of A is 1.0 M and that no B is present at the beginning of the reaction.

Part C
Find the equilibrium concentrations of A and B for a = 1 and b = 2. Assume that the initial concentration of A is 1.0 M and that no B is present at the beginning of the reaction.

2) For the reaction shown here, Kc = 0.513 at 500 K.
N2O4(g)⇌2NO2(g)

If a reaction vessel initially contains an N2O4 concentration of 5.00×10−2 M at 500 K, what are the equilibrium concentrations of N2O4 and NO2 at 500 K?
1
Expert's answer
2020-03-30T10:53:20-0400

1) Part A. A = B

Kc = [B]/[A] = x/(1-x) = 2.7

x is reacted A and produced B

3.7x = 2.7

x = 0.73

[B] = 0.73 M, [A] = 0.27 M

Part B. 2A = 2B

Kc = [B]2/[A]2 = x2/(1-x)2 = 2.7

x reacted A and produced B

x = [B] = 0.62 M, [A] = 0.38 M

Part C. A = 2B

x is reacted A, 2x is produced B

Kc = [B]2/[A] = 4x2/(1-x) = 2.7

x = 0.55, hence, [A] = 0.45 M, [B] = 1.10 M

2) Kc = [NO2]2/[N2O4] = 4x2/(0.05-x) = 0.513

x is reacted N2O4, 2x is produced NO2

x = 0.038

[N2O4] = 0.050 - 0.038 = 0.012 M

[NO2] = 0.076 M